In triangle ABC, AB= 8x=4 and CB= 3x+34 and angles A and C are congruent, find x
Answer by cleomenius(959)

If the angles are congruent, the sides will be the same.
8x + 4 = 3x + 34
5x = 30
x = 6
AB = 8(6) + 4 = 52
CB = 3(6) + 34 = 52
This does check.
Cleomenius.